# From Persona to Personalization: A Survey on Role-Playing Language Agents

### v1

@@ 3.2 RPLA Construction
Role-Playing Language Agents (RPLAs) are primarily developed to simulate intricate personas based on various individual profiles and narratives. These profiles are constructed using diverse persona data, including descriptive narratives, dialogues, historical behaviors, and extensive textual materials such as books (Zhang et al., 2018; Dinan et al., 2020; Shanahan et al., 2023; Wang et al., 2023g; Shao et al., 2023; Xu et al., 2023a; Li et al., 2023f).
The methodologies for building RPLAs typically involve either parametric training (Shao et al., 2023; Wang et al., 2023g; Qin et al., 2024) or nonparametric prompting (Dalvi Mishra et al., 2022; Li et al., 2023a; Zhou et al., 2023a; Gupta et al., 2023; Ma et al., 2023a; Zhao et al., 2023b), both of which may concurrently contribute to the development process. In parametric training, RPLAs undergo pre-training using large-scale raw text, including literary works and encyclopedic entries (Xu et al., 2023a; Gupta et al., 2023), succeeded by supervised fine-tuning focused on persona-specific dialogues (Wang et al., 2023g; Shao et al., 2023). Conversely, nonparametric prompting presents role-playing instructions and examples, based on persona descriptions and dialogue instances (Zhou et al., 2023a; Deshpande et al., 2023; Li et al., 2023a; Shao et al., 2023). Furthermore, modern RPLAs increasingly integrate memory modules to retrieve information from extensive datasets on character traits or past interactions. This development addresses the restricted context capacity of current LLMs (Shao et al., 2023; Mysore et al., 2023; Sun et al., 2024).
@@
In terms of alignment with persona types, parametric learning tends to focus on demographic information and well-known characters, whereas prompting techniques are generally employed for generating fictional personas and highly personalized characters. Current research in RPLA development generally focuses on steering LLMs with demographics (Zhang et al., 2023b; Hong et al., 2023), developing foundation models (Lu et al., 2024; Zhou et al., 2023a), designing agent frameworks (Li et al., 2023a; Wang et al., 2023g) for RPLAs, and crafting persona profiles for specified individuals (Li et al., 2023a; Wang et al., 2023g; Ahn et al., 2023).
@@ 3.3 RPLA Evaluation
For RPLA evaluation, we distinguish the criteria into two primary categories: role-playing capability evaluation for RPLA methodologies, and persona fidelity evaluation for specific personas  (Wang et al., 2024c). The role-playing capabilities of RPLAs are evaluated on their foundation models and construction frameworks, regardless of specific personas. These evaluations concern aspects such as anthropomorphic abilities, attractiveness, and usefulness, which encompass more granular dimensions including conversation ability (Shao et al., 2023), engagement (Zhou et al., 2023a)  [^1] By “engagement”, we mean the state of whether the LLMs are successfully engaged in role-playing activities., persona consistency (Wang et al., 2024c), emotion understanding (Huang et al., 2023a), theory of mind (Kosinski, 2023), and problem-solving ability (Xu et al., 2023a). Persona fidelity, by contrast, concentrates on whether individual RPLAs well replicate the intended personas, including their knowledge (Shao et al., 2023; Li et al., 2023a), linguistic habits (Wang et al., 2023g; Deshpande et al., 2023), personality (Wang et al., 2024c; Huang et al., 2023c), believes (Li et al., 2024a; Wang et al., 2023e), and decision-making (Xu et al., 2024b; Chen et al., 2023a). Overall, current RPLAs have demonstrated promising and improving performance in simulating personas. However, there remain significant gaps between existing RPLAs and fully human-level intelligent agents (Wang et al., 2023g; Chen et al., 2023a).
@@ 4.2 Analysis of Demographics
RPLAs are embedded with intrinsic demographic characteristics, which raises pivotal questions about their ability to role-play specified demographics and the subsequent effects on their behavior. A prevalent approach in demographic role-play involves directly prompting the language agent. For example, if an LLM is prompted with, “You’re a friendly and outgoing individual who thrives on social interactions. Always ready for a good time, you enjoy being the center of attention at parties…” (Jiang et al., 2023a), it adopts the persona of an extroverted character. When tasked with representing distinct demographics, RPLAs demonstrate the capacity to diverge from their inherent traits, manifesting changes in their responses on psychological assessment scales (Jiang et al., 2023a; Serapio-García et al., 2023). This behavioral adaptability highlights the potential of RPLAs in simulating diverse human-like roles and personalities.
@@ 4.3 Application of Demographics
Building upon the capability of single-agent models, which utilize demographic personas to bolster their specialized abilities, assigning demographic personas in multi-agent systems has also emerged as a crucial strategy for enhancing the performance of single-agent systems, i.e., standalone LLMs. By embedding various personas within agents, distinct societal dynamics could be cultivated, leading to improved strategies for cooperative problem-solving and breakthroughs in complex domains such as mathematical modeling (Zhang et al., 2023b; Wang et al., 2023h). A notable implementation of this approach is ChatDev (Qian et al., 2023) and MetaGPT (Hong et al., 2023), frameworks designed specifically for automating software development within a multi-agent conversational platform. In this setup, different agents are assigned specialized roles that collectively contribute to the agile development of software applications. This collaborative model echoes the strategies applied in projects such as OKR-AGENT (Zheng et al., 2023b), where role-specific enhancements within multi-agent architectures have shown to significantly streamline and optimize task execution.
@@
RPLAs have demonstrated remarkable capabilities in simulating nuanced, human-like interactions across various environments. In the realm of gaming, particularly in strategy and role-playing scenarios, RPLAs have shown impressive performance. For example, Chawla et al. (2023) set the agents to be fair or selfish, and shows that selfish agents could contribute not only to their own interests but also to the collective good. Additionally, more elaborate games like Social Deduction Games are particularly illustrative of RPLAs’ capacity to effectively adopt varied roles, as observed in scenarios such as “The Werewolf” (Xu et al., 2023c) and “The Avalon” (Wang et al., 2023d). In diplomacy-focused games such as Cicero, RPLAs have matched or even surpassed human levels of performance et al.(2022)(FAIR), Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al. (FAIR). Similarly, in war simulation games, RPLAs provide valuable insights into the origins of conflicts, enhancing our understanding of complex geopolitical dynamics (Hua et al., 2023). Extending the application of RPLAs beyond gaming environments, these models are also utilized to mimic daily social interactions, thereby narrowing the behavioral gap between artificial agents and humans. This is exemplified in the development of Humanoid Agent frameworks (Wang et al., 2023i), which embody System 1 functionalities—such as basic needs and emotions—to enhance realism and effectiveness in replicating human responses and behaviors. Furthermore, recent findings in multi-agent interaction environments have revealed that diversifying the types of agents, scaling up their number, and increasing interactions, lead to the emergence of unplanned social behaviors. Such behaviors arise spontaneously from discussions among multiple agents, highlighting the potential for complex, dynamic systems within LLM architectures (Gu et al., 2024). This progression from specific gaming applications to broader social simulations illustrates the expanding versatility and depth of RPLAs in understanding and replicating human-like behavior.
@@ 5.1 Definition
Characters are primarily established individuals with their stories widely recognized by the public, including celebrities, historical figures and fictional characters (e.g., Monkey D. Luffy and Hermione Granger). Occasionally, they also include original characters created by individuals (Zhou et al., 2023a). Character RPLAs have recently emerged as a flourishing field of LLM application (e.g., Character.ai), and hence attracted wide research interest as well (Shao et al., 2023; Wang et al., 2023g; 2024c).
For character RPLAs, the essential requirement for effective role-playing is the ability of LLMs to understand characters. Early research has studied character understanding of language models, involving linking descriptions that outline characters’ traits to their roles (i.e., Character Prediction) and personalities (i.e., Personality Understanding): 1) Character prediction mainly focuses on recognizing characters from a provided text. This includes tasks like co-reference resolution (Li et al., 2023c), relationship understanding (Zhao et al., 2024) and character identification (Brahman et al., 2021; Yu et al., 2022; Li et al., 2023c; Zhao et al., 2024). Additionally, some studies investigate if language models can forecast characters’ future actions based on 2) Personality understanding aims to decode character traits from their dialogues and actions, including predicting the depicted traits (Yu et al., 2023) and generating character descriptions (Brahman et al., 2021).
In recent years, LLMs have demonstrated strong capabilities in language understanding and generation, which significantly advanced the development of RPLAs. The research focus in this direction has hence shifted towards applying and promoting LLMs to faithfully reproduce the characters, including their linguistic style (Wang et al., 2023g; Zhou et al., 2023a; Li et al., 2023a; Wang et al., 2023g), knowledge (Li et al., 2023a; Shao et al., 2023; Zhou et al., 2023a; Chen et al., 2023c; Zhao et al., 2023a; Wang et al., 2023g), personality (Shao et al., 2023; Wang et al., 2024c), and even decision-making (Zhao et al., 2023a; Xu et al., 2024b).
Table 1: Datasets for depicting characters. #Character represents the number of characters, with each character having a specific description. #Sample indicates the number of samples. A sample refers to a dialogue or question, and * denotes numbers of multi-turn dialogues. Method describes how samples in the datasets are constructed. Experience Extraction extracts characters’ dialogues or scenes from corresponding origins, while Dialogue Synthesis generates role-playing conversations with advanced LLMs.
Papers #Character #Samples Language Source Method
@@
Character-LLM (Shao et al., 2023) 9 1,900,800 EN Encyclopedia Experience Extraction Dialogue Synthesis
@@
RoleLLM (Wang et al., 2023g) 100 140,726 EN ZH Scripts Experience Extraction Dialogue Synthesis
@@
RoleInteract (Chen et al., 2024) 500 30,800 EN ZH Books Movies Experience Extraction Dialogue Synthesis
LifeChoice (Xu et al., 2024b) 1,401 1,401 EN Literature- Analysis Experience Extraction
InCharacter (Wang et al., 2024c) 32 18,304 EN ZH Personality- Tests Dialogue Synthesis
@@ 5.2 Data for Character RPLAs
Character data is indispensable for the construction of character RPLAs. The data that represents knowledge of these well-established characters can be roughly categorized into two types: 1) Descriptionsdirectly describe the character personas that guide the behaviors of RPLAs. These include various character attributes, such as identity, relationships, and other predetermined attributes. The attributes serve as the knowledge background and are expected to be accurately recalled upon request, such as names and affiliations (Li et al., 2023a; Zhou et al., 2023a; Shao et al., 2023; Wang et al., 2023g; Chen et al., 2023c; Zhao et al., 2023a; Tu et al., 2024; Lu et al., 2024). Additionally, some descriptions further shape the behaviors of RPLAs, such as personality traits (Li et al., 2023a; Wang et al., 2024c). 2) Demonstrations, on the other hand, are representative behaviors of the characters, which reflect their linguistic, cognitive and behavioral patterns  (Li et al., 2023a; Zhou et al., 2023a; Shao et al., 2023; Wang et al., 2023g; Zhao et al., 2023a; Chen et al., 2023c; Tu et al., 2024; Tang et al., 2024; Lu et al., 2024). While RPLAs are not expected to replicate the exact outputs from the demonstration data, they should portray these patterns and generalize to new situations, i.e., producing responses consistent with the demonstrations. Overall, descriptions provide the core and foundational information for RPLAs, while demonstrations, though not mandatory, are also crucial for achieving vividness and fidelity of RPLAs (Wang et al., 2024c).
@@
    Experience Extraction extracts characters’ dialogues or other scenes from original scripts (Li et al., 2023a; Wang et al., 2023g). The extracted scenes faithfully depict the characters. However, understanding and reproducing these scenes for RPLAs may be impractical without more complete background knowledge.
@@
    Dialogue Synthesis synthesizes character conversations using state-of-the-art LLMs to build and augment datasets for character RPLAs. The topics for these conversations could be sourced from corresponding literature (Shao et al., 2023), general task instructions (Wang et al., 2023g), special scenarios such as personality tests (Wang et al., 2024c), and real use cases (Gosling et al., 2023). LLMs could be leveraged to augment the datasets with more role-playing responses by either generating dialogues similar to given ones via in-context learning (Li et al., 2023a), or by role-playing as RPLAs themselves with existing character data to respond to specified topics (Wang et al., 2023g). This process essentially serves as a knowledge distillation of role-playing capabilities from advanced LLMs. However, the quality of synthesized dialogues is limited by teacher LLMs, which often require further filtering (Tu et al., 2024).
@@
In addition, interaction data (mainly conversations) will be continuously produced during the interaction process between RPLAs and individual users, supplementing the original character data. This data further shapes the persona of RPLAs towards users’ individualized preferences, which forks the standard character RPLAs for individual users. This phenomenon concerns both character persona and individualized persona for RPLAs, where studies and analysis remain underexplored.

### v2

@@ 3.2 RPLA Construction
Table 1: An overview of different methods for RPLA construction.
Method Summary
Parametric Training
(Continual) Pre-training Objective: Knowledge injection.
Data: Raw data (books, encyclopedia, etc.).
Advantages: Readily available for well-established demographics and characters; Directly uses the raw data.
Disadvantages: Necessitates training for new personas; May cause catastrophic forgetting.
Supervised Fine-Tuning Objective: Refining role-playing capabilities; Knowledge injection.
Data: Conversation data.
Advantages: Highly effective.
Disadvantages: Necessitates data processing and training for new personas; Potential information loss during data processing.
Reinforcement Learning Objective: Alignment with general or individual users; Improving social reasoning skills.
Data: Conversation data; Preference data;
Advantages: Effective for mitigating harmful content.
Disadvantages: Cost of human preference data; Sparse rewards in multi-agent games.
Nonparametric Prompting
In-context Learning Objective: Knowledge injection; Alignment with individual users.
Data: Raw data; Conversation data.
Advantages: Highly effective; Training-free; Convenient for new personas and personalization; Can incorporate retrieval mechanism for enhanced efficiency.
Disadvantages: May require data processing for new personas; Consumes more tokens and is restricted by context length.
@@
Role-Playing Language Agents (RPLAs) are primarily developed to simulate intricate personas based on various individual profiles and narratives. These profiles are constructed using diverse persona data, including descriptive narratives, dialogues, historical behaviors, and extensive textual materials such as books (Zhang et al., 2018; Dinan et al., 2020; Shanahan et al., 2023; Wang et al., 2024a; Shao et al., 2023; Xu et al., 2023a; Li et al., 2023f).
The methodologies for building RPLAs typically involve either parametric training (Shao et al., 2023; Wang et al., 2024a; Qin et al., 2024) or nonparametric prompting (Dalvi Mishra et al., 2022; Li et al., 2023a; Zhou et al., 2023a; Gupta et al., 2023; Ma et al., 2023a; Zhao et al., 2023b), as summarized in Table 1. These methods may concurrently contribute to the development process.
Parametric training for RPLAs primarily includes pre-training, supervised fine-tuning (SFT) and reinforcement learning (RL). Initially, LLMs for RPLAs are pre-trained on large-scale raw text, including literary works and encyclopedic entries (Xu et al., 2023a; Gupta et al., 2023), equipping them with a broad knowledge of massive demographic and character personas. Subsequently, these LLMs undergo SFT on role-playing datasets (Wang et al., 2024a; Shao et al., 2023), which enhance both their role-playing capabilities and character-specific knowledge. Additionally, RL methods could further refine RPLAs in terms of: 1) Alignment with general users, e.g., improving attractiveness or mitigating harmful content, with preference data from online application users and invited human annotators (RLHF) or synthesized by LLM  (Bai et al., 2022b; Ouyang et al., 2022). 2) Improving social reasoning skills, e.g., in games (Cheng et al., 2024) or goal-driven conversations (Wang et al., 2024b). 3) Alignment with individual users (Shaikh et al., 2024a; Jang et al., 2023).
Conversely, nonparametric prompting provides RPLAs with persona data and role-playing instructions within the context. Prompts for RPLAs primarily consist of persona data that represents the intended personas, including descriptions and demonstrations. Descriptions (or profiles) represent their basic information such as names, backgrounds, experiences, personalities, tones, catchphrases and other attributes (Wang et al., 2024a; Yuan et al., 2024b). Demonstrations illustrate representative behaviors to further align RPLAs with the intended personas, covering dialogues, behaviors, interactions, preferences, stories or other modalities (Li et al., 2023a; Chen et al., 2023c; Dai et al., 2024a). There are several methods for crafting persona data, including: 1) Online Resource Collection, which gathers information from online resources such as Wikipedia, Supersummary and Fandom for widely-known characters (Shao et al., 2023). 2) Automatic Extraction, where LLMs automatically extract persona data such as dialogues from their origins, e.g., books or scripts (Li et al., 2023a). 3) Dialogue Synthesis, which employs advanced LLMs to create and expand role-playing conversation datasets via in-context learning (Li et al., 2023a) or role-playing as the personas (Ran et al., 2024). If provided with corresponding literature for reference (for character personas), this is akin to automatic extraction and the synthesized dialogues are more faithful to the origins. Otherwise, the synthesized data is of limited quality and often necessitates filtering. 4) Human Annotation, which engages human annotators or character fans to summarize persona descriptions or craft high-quality role-playing conversations (Zhou et al., 2023a). Besides, role-playing instructions or requirements could be incorporated to encourage or restrict specific behaviors of RPLAs. Furthermore, modern RPLAs increasingly integrate memory modules to retrieve information from extensive datasets on character traits or past interactions. This development addresses the restricted context capacity of current LLMs (Shao et al., 2023; Mysore et al., 2023; Sun et al., 2024).
@@
In terms of alignment with persona types, parametric learning tends to focus on demographic information and well-known characters, whereas prompting techniques are generally employed for generating fictional personas and highly personalized characters. Current research in RPLA development generally focuses on steering LLMs with demographics (Zhang et al., 2023b; Hong et al., 2023), developing foundation models (Lu et al., 2024; Zhou et al., 2023a), designing agent frameworks (Li et al., 2023a; Wang et al., 2024a) for RPLAs, and crafting persona profiles for specified individuals (Li et al., 2023a; Wang et al., 2024a; Ahn et al., 2023).
@@ 3.3 RPLA Evaluation
For RPLA evaluation, we distinguish the criteria into two primary categories: role-playing capability evaluation for RPLA methodologies, and persona fidelity evaluation for specific personas  (Wang et al., 2024d). The role-playing capabilities of RPLAs are evaluated on their foundation models and construction frameworks, regardless of specific personas. These evaluations concern aspects such as anthropomorphic abilities, attractiveness, and usefulness, which encompass more granular dimensions including conversation ability (Shao et al., 2023), engagement (Zhou et al., 2023a)  [^1] By “engagement”, we mean the state of whether the LLMs are successfully engaged in role-playing activities., persona consistency (Wang et al., 2024d), emotion understanding (Huang et al., 2023a), theory of mind (Kosinski, 2023), and problem-solving ability (Xu et al., 2023a). Persona fidelity, by contrast, concentrates on whether individual RPLAs well replicate the intended personas, including their knowledge (Shao et al., 2023; Li et al., 2023a), linguistic habits (Wang et al., 2024a; Deshpande et al., 2023), personality (Wang et al., 2024d; Huang et al., 2023c), beliefs (Li et al., 2024a; Wang et al., 2023e), and decision-making (Xu et al., 2024c; Chen et al., 2023a). Current methods for evaluation are mainly three-fold: 1) automatic evaluation with ground truth, 2) automatic evaluation without ground truth, 3) multi-choice questions, 4) human-based evaluation. Overall, current RPLAs have demonstrated promising and improving performance in simulating personas. However, there remain significant gaps between existing RPLAs and fully human-level intelligent agents (Wang et al., 2024a; Chen et al., 2023a) that are faithful to the personas, as well as corresponding evaluation methods for more nuanced role-playing.
@@ 4.2 Analysis of Demographics
RPLAs are embedded with intrinsic demographic characteristics, which raises pivotal questions about their ability to role-play specified demographics and the subsequent effects on their behavior. A prevalent approach in demographic role-play involves directly prompting the language agent. For example, if an LLM is prompted with, “You’re a friendly and outgoing individual who thrives on social interactions. Always ready for a good time, you enjoy being the center of attention at parties…” (Jiang et al., 2023a; Xie et al., 2024a), it adopts the persona of an extroverted character. When tasked with representing distinct demographics, RPLAs demonstrate the capacity to diverge from their inherent traits, manifesting changes in their responses on psychological assessment scales (Jiang et al., 2023a; Serapio-García et al., 2023). This behavioral adaptability highlights the potential of RPLAs in simulating diverse human-like roles and personalities.
@@ 4.3 Application of Demographics
Building upon the capability of single-agent models, which utilize demographic personas to bolster their specialized abilities, assigning demographic personas in multi-agent systems has also emerged as a crucial strategy for enhancing the performance of single-agent systems, i.e., standalone LLMs. By embedding various personas within agents, distinct societal dynamics could be cultivated, leading to improved strategies for cooperative problem-solving and breakthroughs in complex domains such as mathematical modeling (Zhang et al., 2023b; Wang et al., 2023g). A notable implementation of this approach is ChatDev (Qian et al., 2023) and MetaGPT (Hong et al., 2023), frameworks designed specifically for automating software development within a multi-agent conversational platform. In this setup, different agents are assigned specialized roles that collectively contribute to the agile development of software applications. This collaborative model echoes the strategies applied in projects such as OKR-AGENT (Zheng et al., 2023b), where role-specific enhancements within multi-agent architectures have shown to significantly streamline and optimize task execution.
@@
RPLAs have demonstrated remarkable capabilities in simulating nuanced, human-like interactions across various environments. In the realm of gaming, particularly in strategy and role-playing scenarios, RPLAs have shown impressive performance. For example, Chawla et al. (2023) set the agents to be fair or selfish, and shows that selfish agents could contribute not only to their own interests but also to the collective good. Additionally, more elaborate games like Social Deduction Games are particularly illustrative of RPLAs’ capacity to effectively adopt varied roles, as observed in scenarios such as “The Werewolf” (Xu et al., 2023c) and “The Avalon” (Wang et al., 2023d). In diplomacy-focused games such as Cicero, RPLAs have matched or even surpassed human levels of performance (FAIR et al., 2022). Similarly, in war simulation games, RPLAs provide valuable insights into the origins of conflicts, enhancing our understanding of complex geopolitical dynamics (Hua et al., 2023). Extending the application of RPLAs beyond gaming environments, these models are also utilized to mimic daily social interactions, thereby narrowing the behavioral gap between artificial agents and humans. This is exemplified in the development of Humanoid Agent frameworks (Wang et al., 2023h), which embody System 1 functionalities—such as basic needs and emotions—to enhance realism and effectiveness in replicating human responses and behaviors. Furthermore, recent findings in multi-agent interaction environments have revealed that diversifying the types of agents, scaling up their number, and increasing interactions, lead to the emergence of unplanned social behaviors. Such behaviors arise spontaneously from discussions among multiple agents, highlighting the potential for complex, dynamic systems within LLM architectures (Gu et al., 2024). This progression from specific gaming applications to broader social simulations illustrates the expanding versatility and depth of RPLAs in understanding and replicating human-like behavior.
@@ 5.1 Definition
Characters are primarily established characters with their stories widely recognized by the public, including celebrities, historical figures and fictional characters (e.g., Monkey D. Luffy and Hermione Granger). Occasionally, they also include original characters created by individuals (Zhou et al., 2023a). Character RPLAs have recently emerged as a flourishing field of LLM application (e.g., Character.ai), and hence attracted wide research interest as well (Shao et al., 2023; Wang et al., 2024a; d).
For character RPLAs, the essential requirement for effective role-playing is the ability of LLMs to understand characters. Early research has studied character understanding of language models, involving linking descriptions that outline characters’ traits to their roles (i.e., Character Prediction) and personalities (i.e., Personality Understanding): 1) Character prediction mainly focuses on recognizing characters from a provided text. This includes tasks like co-reference resolution (Li et al., 2023c), relationship understanding (Zhao et al., 2024) and character identification (Brahman et al., 2021; Yu et al., 2022; Li et al., 2023c; Zhao et al., 2024). Additionally, some studies investigate if language models can forecast characters’ future actions based on; 2) Personality understanding aims to decode character traits from their dialogues and actions, including predicting the depicted traits (Yu et al., 2023) and generating character descriptions (Brahman et al., 2021).
In recent years, LLMs have demonstrated strong capabilities in language understanding and generation, which significantly advanced the development of RPLAs. The research focus in this direction has hence shifted towards applying and promoting LLMs to faithfully reproduce the characters, including their linguistic style (Wang et al., 2024a; Zhou et al., 2023a; Li et al., 2023a; Wang et al., 2024a), knowledge (Li et al., 2023a; Shao et al., 2023; Zhou et al., 2023a; Chen et al., 2023c; Zhao et al., 2023a; Wang et al., 2024a), personality (Shao et al., 2023; Wang et al., 2024d), and even decision-making (Zhao et al., 2023a; Xu et al., 2024c).
Table 2: Datasets for depicting characters. #Char. represents the number of characters, with each character having a specific description. #Samples indicates the number of samples. A sample refers to a dialogue or question, and * denotes the number of multi-turn dialogues. Method describes how samples in the datasets are constructed. Experience Extraction extracts characters’ dialogues or scenes from corresponding origins, while Dialogue Synthesis generates role-playing conversations with advanced LLMs.
Papers #Char. #Samples Lang. Source Method
Established Characters
@@
Character-LLM (Shao et al., 2023) 9 14,300* EN Encyclopedia Experience Extraction Dialogue Synthesis
@@
RoleLLM (Wang et al., 2024a) 100 140,726 EN ZH Scripts Experience Extraction Dialogue Synthesis
@@
SocialBench (Chen et al., 2024) 500 30,800 EN ZH Books Movies Experience Extraction Dialogue Synthesis
TimeChara (Ahn et al., 2024) 14 10,895* EN Books Dialogue Synthesis
LifeChoice (Xu et al., 2024c) 1,401 1,401 EN Literature- Analysis Experience Extraction
MMRole (Dai et al., 2024a) 200 30,800 EN ZH Encyclopedia Dialogue Synthesis
InCharacter (Wang et al., 2024d) 32 18,304 EN ZH Personality- Tests Dialogue Synthesis
Original Characters
PersonaHub (Dai et al., 2024a) 1,000,000 1,000,000 EN - Dialogue Synthesis
@@ 5.2 Data for Character RPLAs
Character data is indispensable for the construction of character RPLAs. The data that represents knowledge of these well-established characters can be roughly categorized into two types: 1) Descriptionsdirectly describe the character personas that guide the behaviors of RPLAs. These include various character attributes, such as identity, relationships, and other predetermined attributes. The attributes serve as the knowledge background and are expected to be accurately recalled upon request, such as names and affiliations (Li et al., 2023a; Zhou et al., 2023a; Shao et al., 2023; Wang et al., 2024a; Chen et al., 2023c; Zhao et al., 2023a; Tu et al., 2024; Lu et al., 2024). Additionally, some descriptions further shape the behaviors of RPLAs, such as personality traits (Li et al., 2023a; Wang et al., 2024d). 2) Demonstrations, on the other hand, are representative behaviors of the characters, which reflect their linguistic, cognitive and behavioral patterns  (Li et al., 2023a; Zhou et al., 2023a; Shao et al., 2023; Wang et al., 2024a; Zhao et al., 2023a; Chen et al., 2023c; Tu et al., 2024; Tang et al., 2024; Lu et al., 2024; Xu et al., 2024b). While RPLAs are not expected to replicate the exact outputs from the demonstration data, they should portray these patterns and generalize to new situations, i.e., producing responses consistent with the demonstrations. Overall, descriptions provide the core and foundational information for RPLAs, while demonstrations, though not mandatory, are also crucial for achieving vividness and fidelity of RPLAs (Wang et al., 2024d).
@@
    Experience Extraction extracts characters’ dialogues or other scenes from original scripts (Li et al., 2023a; Wang et al., 2024a). The extracted scenes faithfully depict the characters. However, understanding and reproducing these scenes for RPLAs may be impractical without more complete background knowledge, making it less suitable to train LLMs with this data.
@@
    Dialogue Synthesis synthesizes character conversations using state-of-the-art LLMs to build and augment datasets for character RPLAs. The topics for these conversations could be sourced from corresponding literature (Shao et al., 2023), general task instructions (Wang et al., 2024a), special scenarios such as personality tests (Wang et al., 2024d), and real use cases (Gosling et al., 2023). LLMs could be leveraged to augment the datasets with more role-playing responses by either generating dialogues similar to given ones via in-context learning (Li et al., 2023a), or by role-playing as RPLAs themselves with existing character data to respond to specified topics (Ran et al., 2024). When referring corresponding literature, this method is close to experience extraction and yields dialogues that are more faithful to the origins. Otherwise, this process essentially serves as a knowledge distillation of role-playing capabilities from advanced LLMs. However, the quality of synthesized dialogu es is limited by teacher LLMs, which often require further filtering (Tu et al., 2024).
@@
In addition, interaction data (mainly conversations) will be continuously produced during the interaction process between RPLAs and individual users, supplementing the original character data. This data further shapes the persona of RPLAs towards users’ individualized preferences, which forks the standard character RPLAs for individual users. This phenomenon concerns both character persona and individualized persona for RPLAs, where studies and analysis remain underexplored. Besides, point-in-time role-playing  (Ahn et al., 2024) presents an area for further study. In practical applications, users may expect RPLAs to role-play characters in a specific time point, e.g., Harry Potter at the age of 5, challenging LLMs to disregard character knowledge beyond the time-point.